# Efficient Automated Circuit Discovery in Transformers using Contextual Decomposition

### v2

@@ Abstract
Automated mechanistic interpretation research has attracted great interest due to its potential to scale explanations of neural network internals to large models. Existing automated circuit discovery work relies on activation patching or its approximations to identify subgraphs in models for specific tasks (circuits). They often suffer from slow runtime, approximation errors, and specific requirements of metrics, such as non-zero gradients. In this work, we introduce contextual decomposition for transformers (CD-T) to build interpretable circuits in large language models. CD-T can produce circuits of arbitrary level of abstraction, and is the first able to produce circuits as fine-grained as attention heads at specific sequence positions efficiently. CD-T is compatible to all transformer types, and requires no training or manually-crafted examples. CD-T consists of a set of mathematical equations to isolate contribution of model features. Through recursively computing contribution of all nodes in a computational graph of a model using CD-T followed by pruning, we are able to reduce circuit discovery runtime from hours to seconds compared to state-of-the-art baselines. On three standard circuit evaluation datasets (indirect object identification, greater-than comparisons, and docstring completion), we demonstrate that CD-T outperforms ACDC and EAP by better recovering the manual circuits with an average of 97% ROC AUC under low runtimes. In addition, we provide evidence that faithfulness of CD-T circuits is not due to random chance by showing our circuits are 80% more faithful than random circuits of up to 60% of the original model size. Finally, we show CD-T circuits are able to perfectly replicate original models’ behavior (faithfulness $=1$) using fewer nodes than the baselines for all tasks. Our results underscore the great promise of CD-T for efficient automated mechanistic interpretability, paving the way for new insights into the workings of large language models.  [^1] CD-T will be released as an open-source python package to benefit the wide mechanistic interpretability research community.
@@ 1 Introduction
Specifically, for evaluation, we use three standard circuit evaluation datasets: indirect object identification (IOI), greater-than comparisons (Greater-than), and docstring completion (Docstring). We compare recovered circuits by CD-T on these datasets with two SOTA baselines, ACDC (Conmy et al., 2023) and EAP (Syed et al., 2023). We demonstrate that CD-T outperforms the baselines by better identifying attention heads in the manual circuits with an average of 97% ROC AUC due to its outstanding ability in recovering negative and supporting heads efficiently. EAP obtains similar runtime as CD-T but performs slightly worse in ROC AUC for Greater-than and IOI. ACDC is the least efficient as it often takes hours to identify a circuit, yet with no better ROC AUC on all tasks compared to other methods. In addition, we demonstrate faithfulness of CD-T circuits is not due to random chance by showing our circuits are 80% more faithful than random circuits of up to 60% of the original model size. Finally, we show CD-T circuits are able to perfectly replicate original models’ behavior (faithfulness $=1$) using fewer nodes  [^2] In our experiments, ”nodes” refers to attention heads or attention heads’ output at a specific sequence. See the general definition in Section 3. than the baselines for all tasks.
@@ 2.2 Circuit discovery methods
Ablation-based methods are essential for identifying key components within models. Through activation patching, prior research on circuit discovery in language models has uncovered subgraphs of attention heads and MLPs that explain model behavior (Wang et al., 2023; Hanna et al., 2023; Heimersheim & Janiak, 2023). However, most of these methods require extensive manual inspection, limiting their application to smaller models (Lindner et al., 2023). To address this, Conmy et al. (2023) introduced Automated Circuit Discovery (ACDC), which scales to models of any size by automatically calculating edge importance between attention heads and MLPs based on model performance on a chosen metric via activation patching. While ACDC is effective, it suffers from slow runtimes and fails to recover certain attention head subsets within the manual circuits. Syed et al. (2023) introduced Edge Attribution Patching (EAP, or ACDC++ as later renamed), which improves ACDC’s efficiency by using a linear approximation of activation patching. However, this reliance on linear approximations can lead to overestimation of edge importance and a weaker correlation with true causal effects. EAP also struggles when the gradient of the performance metric is zero. To resolve these issues, Hanna et al. (2024) proposed Edge Attribution Patching with Integrated Gradients (EAP-IG) for more faithful circuit recovery, although it increases runtime by a constant factor. Another approach trains sparse autoencoders (SAEs) on attention head outputs and uses the SAE-learned features to map attention head contributions to identified circuits. However, SAE-based methods either require carefully crafted human-generated examples (O’Neill & Bui, 2024) or require an impractical amount of compute to train an SAE (He et al., 2024; Marks et al., 2024). Another common limitation of prior methods is the limited level of granularity of the circuits they discover, because most work either doesn’t support further attention head splitting at specific sequence positions (Syed et al., 2023; Hanna et al., 2024; O’Neill & Bui, 2024), or it is fundamentally feasible but requires a prohibitive amount of time to do so (Conmy et al., 2023), which drastically hinders more fine-grained mechanistic interpretability analysis. Our method CD-T addresses the limitations of previous circuit discovery approaches by significantly reducing runtime, using mathematical decomposition equations with no approximations, allowing for arbitrary level of abstraction of circuits, and avoiding both model training and manually crafted examples (i.e. CD-T achieves all the above in an automated manner).
@@ 2.3 Contextual Decomposition in Neural Networks
3 Definitions for circuit discovery algorithms
Before delving into our method, we first formally define a circuit. To maintain consistency in notations with prior work (Shi et al., 2024; Elhage et al., 2021), we view a model as a computational graph $M$, where nodes are activations of the model components in its forward pass (e.g. attention heads and MLPs) and edges are the interactions between those components, and a circuit $C$ is a subgraph of $M$ responsible for certain behavior of interest, such as solving for a specific task. Input embeddings and output logits are treated as input node and output node separately, and are both included in $M$ and $C$.
Naturally, the first step of circuit discovery is to define a task of interest to investigate. This requires specifying a dataset and a task-specific metric to measure the performance of $M$ and $C$. Here we define a task as a tuple $\tau=(D,s)$ of a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ and a score $s\in\mathbb{R}$. The dataset $D$ contains $n$ pairs of inputs $x_{i}\in X$ and outputs $y_{i}\in Y$. The score is a scalar calculated against a task-specific metric, such as the logit difference between IO and S positions in the IOI task. Given an input $x_{i}$, similarly as to how the entire model defines a function $M(x_{i})$ from inputs to logits, we also associate each circuit $C$ with a function $C(x_{i})$, defined by ablating away the effect of all components in $M\backslash C$ (i.e. the components not included in $C$).
@@ 4 Our method: CD-T and mechanistic interpretability
In this section, we first recall the basic operations of transformers (Section 4.1). We then provide technical detail on contextual decomposition and its application to transformers (section 4.2). Finally, we conclude with a circuit discovery algorithm based on CD-T (section 4.3).
@@ 4.2 Contextual decomposition for transformers (CD-T)
Given a decomposition of a vector $x\in\mathbb{R}^{d}$ into relevant and irrelevant constituents $\beta+\gamma=x$ (here we use the word constituents instead of “components” to avoid name collision with the “components of a neural network”, and a module $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$, CD consists of a set of mathematical equations to determine the decomposition of the output of a module $f$ when given $x$ as input: $f(x)\in\mathbb{R}^{k}=\beta_{o}+\gamma_{o}$. The general principle is to find a symbolic expression for $f(x)$ in terms of $\beta$ and $\gamma$, and group the terms in the expression according to whether they are solely a function of $\beta$ or not: the sum of terms solely relying on $\beta$ becomes $\beta_{o}$, and the sum of the remaining terms becomes $\gamma_{o}$. We refer interested readers to Murdoch et al. (2018) and Singh et al. (2019) for further examples of decomposition formulas of various modules. These decomposition rules can be composed through multiple modules (i.e, if $f(\beta_{1},\gamma_{1})=\beta_{2},\gamma_{2}$, and $g(\beta_{2},\gamma_{2})=\beta_{3},\gamma_{3}$, then $g(f(\beta_{1},\gamma_{1}))=\beta_{3},\gamma_{3}$) in order to define decomposition rules for larger computational blocks, including entire neural networks.
@@ 4.3 Automating circuit discovery
Next, we present our algorithm for discovering circuits with CD-T. Although it is possible to perform CD at arbitrary granularity and on arbitrary components of a transformer (e.g., MLPs, query/key/value vectors), in this paper we focus solely on finding circuits consisting of attention heads in a transformer (in particular, we always leave the MLPs fully intact), and our unit of analysis is either the output of an attention head or the output of an attention head at a specific sequence position. For clarity, our method focuses on including or excluding specific nodes from a computational graph, rather than edges; CD-T at once implicitly models both direct and indirect effects of one node to another, and our algorithm does not make a distinction between the two.
In each iteration, we search over potential source nodes to find the ones with highest relevance to a set of target nodes (initialized to the task objective). Once we find the set of nodes which are most relevant to the task, we prune them by some simple heuristic (in our experiments, simply greedily removing the nodes with the least magnitude of impact up to a threshold was often sufficient, or sometimes pruning was not necessary), and designate these most important nodes the target nodes for the next iteration of the algorithm. We halt when the collection of nodes achieves adequate performance (e.g, comparable to the whole network), or has not improved from the previous iteration.
@@
Careful readers may have noticed that this particular metric measures the magnitude of one node’s contribution to another, but not the “direction”, so that nodes which maximize this metric may actually contribute negatively to the task under investigation. We found in our experiments that this does not practically preclude us from finding functioning circuits anyway; it is relatively efficient to greedily prune away nodes by just measuring if doing so improves a circuit’s task performance. On the other hand, this allows us to straightforwardly find nodes which have significant negative contribution to a task metric, such as the “Negative Name Mover Heads” in the IOI task. In principle, it is also possible that different choices of target relevance metric can be developed to adjust the behavior of the circuit discovery algorithm for specific purposes; we do not explore this in the paper but identify it as a possible avenue for further investigation.
Empirically we find, for a fixed target node, the relevance scores of source nodes in different layers to this target node may have different expected magnitudes, due to the numerical effects of propagation through the network. To account for this, we simply normalize the relevance scores by dividing by the average magnitude across scores found in a given layer. Our complete algorithm is described in Algorithm 1, presented is the specific case where we have chosen to decompose our source nodes $s$ so that $\beta_{s}$ is the activation’s deviation from the mean over some distribution.
Algorithm 1 Building a Circuit using CD-T
  Input: data $x$, precomputed mean activations $M$
  Denote by $a_{x}(s)$ the activation of $s$ on input $x$
  Denote by $M(s)$ the precomputed mean activation of $s$ on an appropriate distribution
@@
     for each attention head $s$ upstream of $T$, calculate the relevance of s to $T$ $R(s,T)$
        Run the network from x to s
        $\beta_{s}=a_{x}(s)-M(s)$, $\gamma_{s}=M(s)$
@@
        Use $\beta_{T},\gamma_{T}$ to compute $R(s,T)$ (equation 16)
@@
     Normalize the relevance scores by average magnitude in layer
@@
     Add $S$ to $C$
     Prune nodes from $S$ for which doing so increases the task metric
     If performance of $C$ is satisfactory, return C
     If performance of $C$ is not better than last iteration, return C
     Set the new value of $T$ to $S$
@@ 5.2.1 Runtime and the recovery of manual circuits
Leveraging the manual circuits as a reference standard, we measure how much overlap there is between the recovered circuits and the reference circuits using ROC AUC. Specifically, we measure the ROC AUC by sweeping through a range of thresholds that determine cutoff scores to keep nodes in a circuit. For CD-T, we test by varying the percentile of top nodes to extract in every iteration, in the range of $[90,99]$. At the same time, we also compute the average algorithm runtime across runs when sweeping through the thresholds for each method. Although CD-T supports arbitrary granular representations of the circuit, the runtime is measured by building a “coarse node-level” circuits by not splitting into sequence positions to be comparable with EAP and ACDC because EAP doesn’t support measuring node importance at sequence positions, and ACDC takes an impractical amount of time to achieve sequence-position splitting. To demonstrate CD-T’s capability to identify circuits responsible for underlying algorithm for tasks with great efficiency, we plot the algorithm runtime against the ROC AUC and report the result in Figure 1-Left. From Figure 1-Left we can see CD-T outperforms the baselines by achieving high ROC AUC and low runtime at the same time. EAP obtains similar runtime as CD-T but performs slightly worse in recovering manual circuits for Greater-than and IOI. Even under this coarse nodel-level abstraction, ACDC is the least efficient as it often takes hours to identify a circuit, yet with no better ROC AUC on all tasks compared to other methods. To explain the better ROC AUC performance of CD-T, we find that CD-T is especially good at identifying negative heads and supporting heads, such as the negative name-mover heads and backup name-mover heads in IOI, which other algorithms struggle to do (more discussion in Appendix A).

### v3

@@ Abstract
Automated mechanistic interpretation research has attracted great interest due to its potential to scale explanations of neural network internals to large models. Existing automated circuit discovery work relies on activation patching or its approximations to identify subgraphs in models for specific tasks (circuits). They often suffer from slow runtime, approximation errors, and specific requirements of metrics, such as non-zero gradients. In this work, we introduce contextual decomposition for transformers (CD-T) to build interpretable circuits in large language models. CD-T can produce circuits at any level of abstraction and is the first to efficiently produce circuits as fine-grained as attention heads at specific sequence positions. CD-T is compatible to all transformer types, and requires no training or manually-crafted examples. CD-T consists of a set of mathematical equations to isolate contribution of model features. Through recursively computing contribution of all nodes in a computational graph of a model using CD-T followed by pruning, we are able to reduce circuit discovery runtime from hours to seconds compared to state-of-the-art baselines. On three standard circuit evaluation datasets (indirect object identification, greater-than comparisons, and docstring completion), we demonstrate that CD-T outperforms ACDC and EAP by better recovering the manual circuits with an average of 97% ROC AUC under low runtimes. In addition, we provide evidence that faithfulness of CD-T circuits is not due to random chance by showing our circuits are 80% more faithful than random circuits of up to 60% of the original model size. Finally, we show CD-T circuits are able to perfectly replicate original models’ behavior (faithfulness $=1$) using fewer nodes than the baselines for all tasks. Our results underscore the great promise of CD-T for efficient automated mechanistic interpretability, paving the way for new insights into the workings of large language models. All code for using CD-T and reproducing results is made available on Github.  [^1] https://github.com/adelaidehsu/CD_Circuit
@@ 1 Introduction
Specifically, for evaluation, we use three standard circuit evaluation datasets: indirect object identification (IOI), greater-than comparisons (Greater-than), and docstring completion (Docstring). We compare recovered circuits by CD-T on these datasets with two SOTA baselines, ACDC (Conmy et al., 2023) and EAP (Syed et al., 2023). We demonstrate that CD-T outperforms the baselines by better identifying attention heads in the manual circuits with an average of 97% ROC AUC due to its outstanding ability in recovering negative and supporting heads efficiently. EAP obtains similar runtime as CD-T but performs slightly worse in ROC AUC for Greater-than and IOI. ACDC is the least efficient as it often takes hours to identify a circuit, yet with no better ROC AUC on all tasks compared to other methods. In addition, we demonstrate faithfulness of CD-T circuits is not due to random chance by showing our circuits are 80% more faithful than random circuits of up to 60% of the original model size. Finally, we show CD-T circuits are able to perfectly replicate original models’ behavior (faithfulness $=1$) using fewer nodes  [^2] In our experiments, ”nodes” refers to attention heads or attention heads’ output at a specific sequence. See the general definition in Section 3.3.1. than the baselines for all tasks.
@@ 2.2 Circuit discovery methods
Ablation-based methods are essential for identifying key components within models. Through activation patching, prior research on circuit discovery in language models has uncovered subgraphs of attention heads and MLPs that explain model behavior (Wang et al., 2023; Hanna et al., 2023; Heimersheim & Janiak, 2023). However, most of these methods require multiple iterations of manual inspection and ad-hoc analysis which is specific to the model and task, limiting their application to small numbers of circuits in smaller models (Lindner et al., 2023). To address this, Conmy et al. (2023) introduced Automated Circuit Discovery (ACDC), which scales to models of any size by automatically calculating edge importance between attention heads and MLPs based on model performance on a chosen metric via activation patching. While ACDC is effective, it suffers from slow runtimes and fails to recover certain attention head subsets within the manual circuits. Syed et al. (2023) introduced Edge Attribution Patching (EAP, or ACDC++ as later renamed), which improves ACDC’s efficiency by using a linear approximation of activation patching. However, this reliance on linear approximations can lead to overestimation of edge importance and a weaker correlation with true causal effects. EAP also struggles when the gradient of the performance metric is zero. To resolve these issues, Hanna et al. (2024) proposed Edge Attribution Patching with Integrated Gradients (EAP-IG) for more faithful circuit recovery, although it increases runtime by a constant factor. Another approach trains sparse autoencoders (SAEs) on attention head outputs and uses the SAE-learned features to map attention head contributions to identified circuits. However, SAE-based methods either require carefully crafted human-generated examples (O’Neill & Bui, 2024) or require an impractical amount of compute to train an SAE (He et al., 2024; Marks et al., 2024). Another common limitation of prior methods is the limited level of granularity of the circuits they discover, because most work either doesn’t support further attention head splitting at specific sequence positions (Syed et al., 2023; Hanna et al., 2024; O’Neill & Bui, 2024), or it is fundamentally feasible but requires a prohibitive amount of time to do so (Conmy et al., 2023), which drastically hinders more fine-grained mechanistic interpretability analysis. Our method CD-T addresses the limitations of previous circuit discovery approaches by significantly reducing runtime, using mathematical decomposition equations with no approximations, allowing for arbitrary level of abstraction of circuits, and avoiding both model training and manually crafted examples (i.e. CD-T achieves all the above in an automated manner).
@@ 3 Our method: CD-T and mechanistic interpretability
In this section, we first recall the basic operations of transformers (Section 3.1). We then provide technical details on contextual decomposition and its application to transformers (section 3.2). Finally, we conclude with a circuit discovery algorithm based on CD-T (section 3.3).
@@ 3.2 Contextual decomposition for transformers (CD-T)
Given a decomposition of a vector $x\in\mathbb{R}^{d}$ into relevant and irrelevant constituents $\beta+\gamma=x$ (here we use the word constituents instead of “components” to avoid name collision with the “components of a neural network”), and a module $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$, CD consists of a set of mathematical equations to determine the decomposition of the output of a module $f$ when given $x$ as input: $f(x)\in\mathbb{R}^{k}=\beta_{o}+\gamma_{o}$. The general principle is to find a symbolic expression for $f(x)$ in terms of $\beta$ and $\gamma$, and group the terms in the expression according to whether they are solely a function of $\beta$ or not: the sum of terms solely relying on $\beta$ becomes $\beta_{o}$, and the sum of the remaining terms becomes $\gamma_{o}$. We refer interested readers to Murdoch et al. (2018) and Singh et al. (2019) for further examples of decomposition formulas of various modules. These decomposition rules can be composed through multiple modules (i.e, if $f(\beta_{1},\gamma_{1})=\beta_{2},\gamma_{2}$, and $g(\beta_{2},\gamma_{2})=\beta_{3},\gamma_{3}$, then $g(f(\beta_{1},\gamma_{1}))=\beta_{3},\gamma_{3}$) in order to define decomposition rules for larger computational blocks, including entire neural networks.
@@ 3.3 Automating circuit discovery
3.3.1 Definitions of circuit discovery
Before delving into our method, we first formally define a circuit. To maintain consistency in notations with prior work (Shi et al., 2024; Elhage et al., 2021), we view a model as a computational graph $M$, where nodes are activations of the model components in its forward pass (e.g. attention heads and MLPs) and edges are the interactions between those components, and a circuit $C$ is a subgraph of $M$ responsible for certain behavior of interest. Naturally, the first step of circuit discovery is to define a task of interest to investigate. This requires specifying a dataset and a task-specific metric to measure the performance of $M$ and $C$, typically, by maximization of the task metric. Given an input $x$, similarly as to how the entire model defines a function $M(x)$ from inputs to logits, we also associate each circuit $C$ with a function $C(x)$, defined by ablating away the effect of all components in $M\backslash C$ (i.e. the components not included in $C$). Circuit discovery may be performed on a single input example or, likelier, the result may be averaged over multiple input examples to reduce variance. Typically we find averaging across 10-100 samples, depending on the tasks being studied, can yield stable performance. Detailed information about which models are studied, input sample counts, and how the task metrics are defined is found in Appendix A. Finally, note that our method finds the circuit of interest but does not automatically generate an interpretation of it, however, known interpretability techniques (Wang et al., 2023) can be used to do so.
3.3.2 Method
Here we present our algorithm for discovering circuits with CD-T. Although it is possible to perform CD at arbitrary granularity and on arbitrary components of a transformer (e.g., MLPs, query/key/value vectors), in this paper we focus solely on finding circuits consisting of attention heads in a transformer to be comparable to prior work, and our unit of analysis is either the output of an attention head or the output of an attention head at a specific sequence position. For clarity, our method focuses on including or excluding specific nodes from a computational graph, rather than edges; CD-T at once implicitly models both direct and indirect effects of one node to another, and our algorithm does not make a distinction between the two.
In each iteration, we search over potential source nodes to find the ones with highest relevance to a set of target nodes (initialized to the task objective). Once we find the set of nodes which are most relevant to the task, we prune them by some simple heuristic (in our experiments, simply greedily removing the nodes with the least magnitude of impact was often sufficient, or sometimes pruning was not necessary), and designate these most important nodes the target nodes for the next iteration of the algorithm. We halt when the collection of nodes achieves adequate performance (e.g, comparable to the whole network), or has not improved from the previous iteration.
@@ 3.3.2 Method
Careful readers may have noticed that this particular metric measures the magnitude of one node’s contribution to another, but not the “direction”, so that nodes which maximize this metric may actually contribute negatively to the task under investigation. We found in our experiments that this does not preclude us from finding functioning circuits. On the contrary, this allows us to straightforwardly find nodes which have significant negative contribution to a task metric, such as the “Negative Name Mover Heads” in the IOI task. In principle, it is possible that different choices of target relevance metric can be developed to adjust the behavior of the circuit discovery algorithm to avoid (or accentuate) this behavior; we do not explore this in the paper but identify it as a possible avenue for further investigation. Our complete algorithm is described in Algorithm 1, presented in the specific case where we have chosen to decompose our source nodes $s$ so that $\beta_{s}$ is the activation’s deviation from the mean over some distribution. More details on the heuristics used and a complexity analysis of the algorithm is found in Appendix B.
Algorithm 1 Building a circuit using CD-T for a specified task
  Input: datapoint $x$, precomputed mean activations $\mu$
  Denote $a_{x}(s)$ as the activation of $s$ on input $x$
  Denote $\mu(s)$ as the precomputed mean activations of $s$ on an appropriate distribution
@@
     for each attention head $s$ upstream of $T$
        Run the model from $x$ to $s$, and compute $\beta_{s}=a_{x}(s)-\mu(s)$, $\gamma_{s}=\mu(s)$
@@
        Use $\beta_{T},\gamma_{T}$ to compute the relevance of s to $T$, $R(s,T)$ using equation 16
@@
     $C\leftarrow C\cup S$# Refinement: Greedily prune away the unnecessary nodes}
     repeat
        for each node $n$ in $C$
           $C^{{}^{\prime}}\leftarrow C\backslash n$
           If $C^{{}^{\prime}}(x)>C(x)$: $C\leftarrow C^{{}^{\prime}}$
        end for
     until $C$ has not changed since the last iteration
     If $|C(x)-M(x)|<\epsilon$, return $C$# Circuit performs close to the full model}
     If $C(x)$ is not greater than last iteration, return $C$
     $T\leftarrow S$
@@ 4.2.1 Runtime and the recovery of manual circuits
Leveraging the manually discovered circuits as a reference standard, we measure how much overlap there is between the recovered circuits and the reference circuits using ROC AUC. Specifically, we measure the ROC AUC by sweeping through a range of thresholds that determine cutoff scores to keep nodes in a circuit. For CD-T, we test by varying the percentile of top nodes to extract in every iteration, in the range of $[90,99]$. At the same time, we also compute the average algorithm runtime across runs when sweeping through the thresholds for each method. Although CD-T supports arbitrary granular representations of the circuit, the runtime is measured by building a “coarse node-level” circuits by not splitting into sequence positions to be comparable with EAP and ACDC because EAP doesn’t support measuring node importance at sequence positions, and ACDC takes an impractical amount of time to achieve sequence-position splitting. To demonstrate CD-T’s capability to identify circuits responsible for underlying algorithm for tasks with great efficiency, we plot the algorithm runtime against the ROC AUC and report the result in Figure 1-Left. From Figure 1-Left we can see CD-T outperforms the baselines by achieving high ROC AUC and low runtime at the same time. EAP obtains similar runtime as CD-T but performs slightly worse in recovering manual circuits for Greater-than and IOI. Even under this coarse nodel-level abstraction, ACDC is the least efficient as it often takes hours to identify a circuit, yet with no better ROC AUC on all tasks compared to other methods. To explain the better ROC AUC performance of CD-T, we find that CD-T is especially good at identifying negative heads and supporting heads, such as the negative name-mover heads and backup name-mover heads in IOI, which other algorithms struggle to do (more discussion in Appendix A).